OpenStudy (niahloves):

PLlease help!! medal and fan Which equation can be used to find the answer? Sharifa spends $51 on the first day of her vacation. This leaves her with $127 to spend during the rest of her vacation. How much money did Sharifa bring to spend during her vacation? A. t = 127 · 2 t = 254 $254 B. t – 51 = 127 t = 178 $178 C. t + 51 = 127 t = 76 $76 D. t = 127 ÷ 2 t = 63.5 $63.50

OpenStudy (anonymous):

B

OpenStudy (niahloves):

Which equation can be used to find the answer? Christina climbs to the top of Mount Graham in Arizona. On the first day, she climbs to an elevation of 1029 ft. The elevation of Mount Graham is 10,724 ft. How many total feet of Mount Graham’s elevation does Christina climb on the other days? A. e + 1029 = 10,724 e = 9695 9695 ft B. e = 1029 · 2 e = 2058 2058 ft C. e = 10,724 ÷ 2 e = 5362 5362 ft D. e – 1029 = 10,724 e = 11,753 11,753 ft @Firez

OpenStudy (niahloves):

@Conqueror

OpenStudy (anonymous):

It's A